# Center Biophysics and Biochemistry, (IVIC), Caracas, Venezuela

### v2

@@ Abstract
Mandelbrot [137, 92] has pointed out, that the fractal dimension is related to previous concepts associated with, for example, the flow of rivers, which would be an interesting natural phenomenon per se. This would be the case, of the flow of the Nile River. Harold Edwin Hurst (1880 – 1978) a British hydrologist [Sutcliffe1979, Sutcliffe1999, 158, Sutcliffe2016] who was dedicated to measuring the level of the River since 1906 for 62 years. Hurst published his first work in 1951, one of the most influential and highly cited works in scientific hydrology, [105], he was then 71 years old. The work of 1951 is followed by others in 1956 [106, 108, 107].
Figure 11: Credits of the figure in Recognition [107].
Hurst [107] has a record of the flow of the Nile, yours since 1906 ($\approxeq 44$ years) preceded by about $\geqslant 2000$ [107, Table 2 and Figure 1] years of historical record. If we assume that the flow of the Nile enters a reservoir of indefinite capacity, from which there is a constant efflux equal to the annual average of discharge (evaporation, irrigation, consumption, etc.). The storage that would be required to make the discharge flow of each year is obtained by computing the continuous sums of the annual deviations from the mean. Then the range $R$ between the maximum and minimum of these continuous sums is either: (a) the maximum storage when there is no deficit, (b) the maximum accumulated deficit when there has never been maximum storage or (c) its sum when there is storage and deficit.
@@
The storage required for the annual discharge to flow is obtained by adding the excess water to the average. Then the range $\boldsymbol{{R}}$ from maximum to minimum of the continuous sums is either ($a$) the average accumulated storage when there is never a deficit, ($b$) the maximum accumulated deficit when there is never a water accumulation, or ($c$) the sum when there is accumulated and deficit water. In Figure 11, $\boldsymbol{{R=(P\text{---}C)+(D\text{---}Q)}}$ [107, pg. 14].
This is shown in Figure 11 [107, Figure 1] where $\boldsymbol{{X\in(0,X_{max})}}$ is the time axis and $\boldsymbol{{Y\in(-Y_{min},Y_{max})}}$ is the axis of accumulated departures from the mean $\boldsymbol{{M=\overline{Y}}}$, of $\boldsymbol{{Y}}$. The curve expressing this transitions, in sequence, through the points $\boldsymbol{{0\rightarrow P\rightarrow Q\rightarrow E}}$, and $\boldsymbol{{Q\rightarrow E}}$. $\boldsymbol{{N}}$ is the number of year of observation [107, pg. 14]. During the period $\boldsymbol{{0\rightarrow C}}$ the deviations from the mean are positive and the damn storage or river level of the interval, increases from $\boldsymbol{{0\rightarrow P}}$. Then $\boldsymbol{{C\rightarrow D}}$ period, represents a reduction in water storage in period $\boldsymbol{{P\longrightarrow Q}}$, which determines that the curve reaches a negative level with respect to the mean storage and then returns to the mean value $\boldsymbol{{Q\rightarrow 0}}$ in the $\boldsymbol{{D\rightarrow E}}$, period. $\boldsymbol{{P\rightarrow C+D\rightarrow Q=R}}$, and this amount of storage would have allowed the average discharge of the period to have been maintained during it. The point $F$ whose ordinate is $\boldsymbol{{E\text{---}F=\left[-N(M-B)\right]}}$. Hurst notation is, generally speaking awkward, $\boldsymbol{{B}}$ is water draft which is $\boldsymbol{{B<M}}$ [107, pg. 19]; $\boldsymbol{{F}}$ is thus an indicator of water overdraw.
@@
Hurst [107] usually assumes that the flow distribution of rivers approximates a Normal or Gaussian curve, and that this is usually true, but that it is only part of the description of the phenomenon, since there is also a tendency to exist for high and low years that are grouped together. “Theoretical research shows that, if individual years were completely independent of each other and each year’s downloads were completely independent of each other and downloads, the most likely value of $\boldsymbol{{R|\sigma}}$ would be given by[107]
$\tfrac{R}{\sigma}=\sqrt{\tfrac{\pi}{2}N}=(1.25\ldots)\cdot\sqrt{N},$ (34)
where, as said, $\boldsymbol{{N}}$ is the number of years and $\boldsymbol{{\sigma}}$ is the standard deviation of downloads for the period considered. Experiments with random events such as flipping coins agree with this equation.
Records of discharges from a number of rivers were examined and $R$ was calculated for as many of them as were available. Unfortunately, no records were then found of discharges from a river covering more than 70 years, and therefore $\boldsymbol{{R}}$ was computed for a number of rainfall records of which several spanned more than 150 years. To these was added some records of various river levels, temperatures, and pressures. A common feature of all records was that their frequency distributions, ignoring the order in which they occurred, were of the rounded type to approximate the curve of the Gaussian normal distribution. When $\boldsymbol{{R}}$ was plotted as a function of $\boldsymbol{{N}}$ the equation produced an elongated group of points
@@
In this group there was nothing to distinguish one type of phenomenon from another, but it was clear that that $\boldsymbol{{R}}$ was increasing faster than was the case with random events. This was attributed to the tendency of natural phenomena to have runs when the values as a whole higher values high and others when they were low. Due to the dispersion of points whose length of available records was not large enough to decide whether $\boldsymbol{{\tfrac{R}{\sigma}}}$ for natural phenomena was better represented by the square root or by some other function of $\boldsymbol{{N}}$.
@@
In the attempt to resolve this point longer records of rainfall, temperatures, pressure, and lake levels continued to be computed for longer natural periods of natural phenomena. as a result the analysis was extended to the records of the Nilometer of Roda (Cairo), which went back, with discontinuities, to the year 640 BC, the thickness that produced records up to 900 years, and clay tablets from which 4,000 years of record were available.
@@
In total, 75 different phenomena were used. In the case of three rings results from 4 different locations were considered separately, and the findings used were separate, and the findings were used for each locality are the means of a group of approximately 10 trees (“trees” in the original). A record was divided into periods and for each of which the $\boldsymbol{{\tfrac{R}{\sigma}}}$ was compared. For example, with 120 years recorded it can be calculated for 3 periods of 40 years, two covering periods of 80, and a full period of 120 years. In general $R$ was computed for periods of less than 30 years. en total 690 values of $\boldsymbol{{\tfrac{R}{\sigma}}}$ were used. A preliminary long-term examination of the data showed that $\tfrac{R}{\sigma}$ increased faster than $\sqrt{N}$ and less rapidly than and $\boldsymbol{{N}}$. To find the form of the relationship the statistics were divided into sets containing similar phenomena, and the sets were again divided into groups, each group consisting of a small number of values of $\boldsymbol{{\tfrac{R}{\sigma}}}$ with approximately the same value of $\boldsymbol{{N}}$. Hurst [107] tabulates these values in his Table 1 as means of $\boldsymbol{{\tfrac{R}{\sigma}}}$ together with $\boldsymbol{{N}}$ and its logarithms. Figure 2 of Hurst [107] presents the $\boldsymbol{{\tfrac{R}{\sigma}}}$ plotted against $\boldsymbol{{\log(N)}}$ to produce 7 good linear regressions.
@@
A striking point of the discussion that Hurst in all his works, is that he uses concepts such as media, deviation from the mean, standard deviation ($\boldsymbol{{\sigma}}$) or “Gaussian” without any explanation or justification. It seems to mean data distributed with a pdf such as $\boldsymbol{{\text{{{N}}}\left\{\mu,\sigma^{2}\right\}}}$, which should be distributed symmetrically around d $\boldsymbol{{\mu}}$, however that symmetry does not exist in our Figure 11 [107, Original Figure 1] or in any equivalent figure in any Hurst work. As stated above, if the median and 95% CI are calculated for the 11 values of $\boldsymbol{{K}}$ in the in the [107, Table 1], there is an asymmetry between the medians and their 95% CI for the values of $K$, even though we do not place the values here because the samples become small, the asymmetry also exists for the sets of $\boldsymbol{{K}}$ separating the data of $\boldsymbol{{N=114}}$ and $N=25$ from [107, Table 1]. This probably ignorance of Hurst and the tendency until the first half of the twentieth century to consider all Gaussian data and call them “normal”, in the sense of being the “norm”.
@@
Hurst´s [107, Figure 2] shows that there is a linear relationship between $\boldsymbol{{log(\tfrac{R}{\sigma})}}$ and $\boldsymbol{{\log\left(N\right)}}$ for all sets in which there are enough groups. The equations of these lines are (for $\boldsymbol{{1.8\lessapprox\log(N)\lessapprox 3}}$ [107, in Figure 2])
$\log\left(\tfrac{R}{\sigma}\right)=K\log\left(\tfrac{N}{2}\right).$ (35)
@@
$\boldsymbol{{\left(\tfrac{R}{\sigma}\right)=R_{\sigma}}}$ is called the rescaled range of fluctuations, in this case, of the Nile [158]. The value of $\boldsymbol{{K}}$ in Table 2 of Hurst [107] is positive and (calculated by me) 0.70(0.69 – $\boldsymbol{{0.70\;(0.69\text{ -- }0.72)}}$ (median and 95% confidence interval, calculated according to Hodges and Lehmann [98]), Very little dispersed, although perhaps somewhat skewed upwards. It is clear from the figures that for each set of phenomena a straight line fits the data well, and the remarkable fact that $\boldsymbol{{K}}$, the slope of the line, varies little from one set to another. A summary of the mean of the $K$ values obtained for the 690 separate values that Hurst computed, which is equivalent to $\boldsymbol{{\approx\tfrac{1,\bar{3}}{month}}}$ of the values that Hurst measured in years in which he devoted himself to that. The fraction $\boldsymbol{{\approx\tfrac{1,\bar{3}}{month}}}$ perishes little, but probably reflects what would be expected of a great river: that its changes are slow. The number of data used $\boldsymbol{{\approx\tfrac{1}{month}}}$, is curious since some contemporary information suggests that the level of the Nile is measured daily [4].
@@
So far we have followed as closely as we could the data from Hurst [107]. Here we must point out however that Eq. (35) can be rewritten in general form as
@@
which suggests that straight lines are obtained by Brando logarithms of any base. From here on we will join most authors calling $\boldsymbol{{K_{b}}}$ Hurst’s coefficient and denote it as $\boldsymbol{{H}}$. An extremely important particular case is
$\begin{split}\ln\left(\dfrac{R}{\sigma}\right)&=\ln\left(R_{\sigma}\right)=H\ln\left(\dfrac{N}{2}\right)\\ \boldsymbol{{\text{\wasytherefore{\quad}}R}}&\boldsymbol{{=\left(\dfrac{N\sigma}{2}\right)\text{{{e}}}^{H}\implies R_{\sigma}=\left(\dfrac{N}{2}\right)\text{{{e}}}^{H}\qquad\boldsymbol{\square}}}\end{split}$ (37)
Hurst [107, p. 19] considers the case where the annual water demand is less than the average river flow, but not less than the minimum annual flow recorded. In this case you have an approximate, $\boldsymbol{{B}}$, equal to what is required and less than the average, $\boldsymbol{{M}}$, and we need to know the largest accumulated deficit, $\boldsymbol{{S}}$, which must be covered with stored water. This can be determined for any particular record of the cumulative deviation curve.
Referring to the Figure 11, $\boldsymbol{{0\rightarrow X}}$ is the abscissa, which represents the value of the mean, with respect to which the other values of the figure are graphed. If we now produce an estimate $\boldsymbol{{B}}$ less than the average storage (axis $\boldsymbol{{0\rightarrow X}}$) it will be increased each year in the magnitude $\boldsymbol{{M-B}}$ above what was the estimated $\boldsymbol{{M}}$. This can be determined from the original curve of cumulative deviations by drawing an axis $\boldsymbol{{0\rightarrow X}}$ passing through the point $F$ whose ordinate is $\boldsymbol{{-N(M-B)}}$. The curve ordinates referred to the $\boldsymbol{{0axis\rightarrow X_{1}}}$ show storage changes with water withdrawals $\boldsymbol{{B}}$. Storage drops from $\boldsymbol{{P}}$ to $Q$, $\boldsymbol{{P\rightarrow c+d\rightarrow Q}}$. This is the amount of storage $S$ that would have been necessary to meet the demand $\boldsymbol{{B}}$.
$\boldsymbol{{S}}$ was determined for various water demands for each of $38$ worth of phenomena taken from the kinds of river discharges, rainfall, evaporation, use, and temperature that determines evaporation. The results for each class were grouped according to extraction, and the means of these groups were plotted. Evaporation is currently an important factor determining the salinization of Lake Nasser [7]. Two relationships fix the results really well. The adjustment is (this strange equation is Hursy´s verbatim [107, Eq. (3)]):
@@
From there Hurst [107, Eq. (4)]) somehow gets
@@
The average value of $\boldsymbol{{N}}$ from which these results are deducted is $96$. It will be seen that in the vicinity of the adjustment, on the range of observations, there is no significant difference between one type of relationship and the other. At some future time it may decide one of the types has some potential theoretical theoretical justification. If the figures are examined you will notice values of these kinds of phenomena are indistinguishable.
Hurst data on thr Nile river is indeed priceless, but as seen in Eqs. (38) and (39), his algebra is sometimes strange strange. Another source of uncertainty in Hurst´s data analysis is his unproven data Gaussianity, which is important to relate his exponent (now called H) with D, the fractal dimension of Nile´s water level fluctuations [141, 68, 69, Sutcliffe2016].
5.3 Validity of the relationship between the Hurst´s coefficient and the fractal dimension.
The origin of the association between the two parameters was initially proposed by Mandelbrot and Wallis [144], with Mandelbrot’s success in finding fractal systems at the most diverse sites there seemed to be evidence that between the Hurst exponent and the fractal dimension, $\Phi$, of two-dimensional was a simple relationship such as
@@
[137, 92, Sevcik1998a, Sevcik2010]. Thanks to Mandelbrot’s success, fractal analysis has been applied to time series, profiles, and natural surfaces in almost every scientific discipline. More recent analyses [141, 68, 69, Sutcliffe2016] indicate that the relationship between fractal dimension and Hurst’s coefficient is actually more complex than initially thought, and is not described by Eq. (40).
Because of this, several authors have developed independent and (more?) exact methods to evaluate the Hurst coefficient [69, 66, Sanchez2015] which do not presuppose the Gaussianityd of the data that Hurst made and even consider the effect of the data ‘pathological’ distributed with FDP de Cauchy [48, 167, Walck1996, Wolfram2003]. The discussion of “pathological distributions” like Cauchy’s is neither trivial nor irrelevant. As we discussed in Chapter 6, these distributions are not Gaussian, and have neither mean nor variance. The data that Hurst discusses are deviations from the level of the Nile from its mean, estimated $\boldsymbol{{\bar{x}}}$, this is a random variable, whose mean is
@@
${\text{var}\left[{\Delta x_{i}}\right]}=\dfrac{n+1}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}.$ (42)
If $\boldsymbol{{x_{i}}}$ were Gaussian, then $\Delta x_{i}$ would have PDF and pdf of Cauchy and no operation with its apparent mean and/or apparent variance would make sense. The Eqs. (41) and (42) can be calculated but are meaningless for a Cauchy distribution without central moments [77, 79, 80, 81, Theiler1990, 78]. If $\boldsymbol{{x_{i}}}$ is not Gaussian the problem is similar, an unknown pathological random variable whose mean and variance estimated from a sample, are meaningless, because the pdf of count come has no mean or defined variance. All sampling theory makes sense if and only if, the parameters that are estimated from the sample are approximations of the parameters of the population we sample.
@@
The idea that certain systems, numerically apparently very simple, could not have predictable results until at short times appeared when Lorenz [131] tried to solve a model of climate. The Lorenz model focused on three simultaneous equations (with slightly modified notation):

### v3

@@ Abstract
Mandelbrot [139, 94] has pointed out, that the fractal dimension is related to previous concepts associated with, for example, the flow of rivers, which would be an interesting natural phenomenon per se. This would be the case, of the flow of the Nile River. Harold Edwin Hurst (1880 – 1978) a British hydrologist [Sutcliffe1979, Sutcliffe1999, 160, Sutcliffe2016] who was dedicated to measuring the level of the Nile river since 1906, for 62 years. Hurst published his first paper in 1951, one of the most influential and highly cited works in scientific hydrology, [107], he was then 71 years old. Hurst’s nomenclature is not easy to follow, since almost all of his variables are single letters hinting nothing about their meaning. The work of 1951 is followed by other papers in 1956 [108, 110, 109].
Figure 11: Credits of the figure in Recognitions [109].
Hurst [109] recorded the level of the Nile, since 1906 ($\approxeq 44$ years) preceded by about $\geqslant 2000$ [109, Table 2 and Figure 1] years of historical records. If we assume that the flow of the Nile enters a reservoir of indefinite capacity, from which there is a constant efflux equal to the annual average of discharge (evaporation, irrigation, consumption, etc.). The storage that would be required to make the discharge flow of each year is obtained by computing the continuous sums of the annual deviations from the mean. Then the range $R$ between the maximum and minimum of these continuous sums is either: (a) the maximum storage when there is no deficit, (b) the maximum accumulated deficit when there has never been maximum storage or (c) its sum when there is storage and deficit.
@@
The storage required for the annual discharge to flow is obtained by adding the excess water to the average. Then the range $\boldsymbol{{R}}$ from maximum to minimum of the continuous sums is either: (a), the average accumulated storage when there is never a deficit; (b), the maximum accumulated deficit when there is never a water accumulation; or (c), the sum when there is accumulated and deficit water. In Figure 11 (or [109, pg. 14]),
$\boldsymbol{{R=(P\text{---}C)+(D\text{---}Q)}}.$ (35)
This is shown in Figure 11 [109, Figure 1] where $\boldsymbol{{X\in(0,X_{max})}}$ is the time axis and $\boldsymbol{{Y\in(-Y_{min},Y_{max})}}$ is the axis of accumulated departures from the mean $\boldsymbol{{M=\overline{Y}}}$, of $\boldsymbol{{Y}}$. The curve expressing this transitions, in sequence, through the points $\boldsymbol{{0\rightarrow P\rightarrow Q\rightarrow E}}$, and $\boldsymbol{{Q\rightarrow E}}$ 666$\rightarrow$ is used as “approaches” or “tends to”.. $\boldsymbol{{N}}$ is the number of year of observation [109, pg. 14]. During the period $\boldsymbol{{0\rightarrow C}}$ the deviations from the mean are positive and the damn storage or river level of the interval, increases from $\boldsymbol{{0\rightarrow P}}$. Then $\boldsymbol{{C\rightarrow D}}$ period, represents a reduction in water storage in period $\boldsymbol{{P\longrightarrow Q}}$, which determines that the curve reaches a negative level with respect to the mean storage and then returns to the mean value $\boldsymbol{{Q\rightarrow 0}}$ in the $\boldsymbol{{D\rightarrow E}}$, period. $\boldsymbol{{P\rightarrow C+D\rightarrow Q=R}}$, and this amount of storage would have allowed the average discharge of the period to have been maintained during it. The point $F$ whose ordinate is $\boldsymbol{{E\text{---}F=\left[-N(M-B)\right]}}$. Hurst notation is, generally speaking awkward, $\boldsymbol{{B}}$ is water draft which is $\boldsymbol{{B<M}}$ [109, pg. 19]; $\boldsymbol{{F}}$ is thus an indicator of water overdraw.
@@
[…]
$R\tfrac{R}{\sigma}=\sqrt{\tfrac{\pi}{2}N}=(1.25\ldots)\cdot\sqrt{N},$ (36)
where, as said, $\boldsymbol{{N}}$ is the number of years of observation and $\boldsymbol{{\sigma}}$ is the standard deviation of dischargesfor the period considered. Experiments with random events such as flipping coins agree with this equation.
Records of discharges from a number of rivers were examined and $R$ was calculated for as many of them as were available. Unfortunately, no records were then found of discharges from a river covering more than 70 years, and therefore $\boldsymbol{{R}}$ was computed for a number of rainfall records of which several spanned more than 150 years. To these were added some records of various river levels, temperatures, and pressures. A common feature of all records was that their frequency distributions, ignoring the order in which they occurred, were of the rounded type to approximate the curve of the Gaussian normal distribution. When $\boldsymbol{{R}}$ was plotted as a function of $\boldsymbol{{N}}$ the equation produced an elongated group of points
@@
In this group there was nothing to distinguish one type of phenomenon from another, but it was clear that that $\boldsymbol{{R}}$ was increasing faster than was the case with random events. This was attributed to the tendency of natural phenomena to have runs when the values as a whole are higher values and others were low. Due to the dispersion of points whose length of available records was not large enough to decide whether $\boldsymbol{{\tfrac{R}{\sigma}}}$ for natural phenomena was better represented by the square root or by some other function of $\boldsymbol{{N}}$.
@@
In the attempt to resolve this point longer records of rainfall, temperatures, pressure, and lake levels continued to be computed for longer natural periods of natural phenomena. as a result the analysis was extended to the records of the Nilometer of Roda (Cairo, Egypt), which went back, with discontinuities, to the year 640 BC, the thickness that produced records up to 900 years, and clay tablets from which 4,000 years of record were available.
@@
In total, 75 different phenomena were used. In the case of three rings results from 4 different locations were considered separately, and the findings were separate, and the findings were used for each locality are the means of a group of approximately 10 trees . A record was divided into periods and for each of them the $\boldsymbol{{\tfrac{R}{\sigma}}}$ was compared. For example, with 120 years recorded it can be calculated for 3 periods of 40 years, two covering periods of 80, and a full period of 120 years. In general $R$ was computed for periods of less than 30 years, in total 690 values of $\boldsymbol{{\tfrac{R}{\sigma}}}$ were used. A preliminary long-term examination of the data showed that $\tfrac{R}{\sigma}$ increased faster than $\sqrt{N}$ and less rapidly than and $\boldsymbol{{N}}$. To find the form of the relationship the statistics were divided into sets containing similar phenomena, and the sets were again divided into groups, each group consisting of a small number of values of $\boldsymbol{{\tfrac{R}{\sigma}}}$ with approximately the same value of $\boldsymbol{{N}}$. Hurst [109] tabulates these values in his Table 1 as means of $\boldsymbol{{\tfrac{R}{\sigma}}}$ together with $\boldsymbol{{N}}$ and its logarithms. Figure 2 of Hurst [109] presents the $\boldsymbol{{\tfrac{R}{\sigma}}}$ plotted against $\boldsymbol{{\log(N)}}$ to produce 7 good linear regressions.
@@
A striking point of the discussion that Hurst in all his works, is that he uses concepts such as media, deviation from the mean, standard deviation ($\boldsymbol{{\sigma}}$) or “Gaussian” without any explanation or justification. It seems to mean data distributed with a pdf such as $\boldsymbol{{\text{{{N}}}\left\{\mu,\sigma^{2}\right\}}}$, which should be distributed symmetrically around d $\boldsymbol{{\mu}}$, however that symmetry does not exist in our Figure 11 [109, Original Figure 1] or in any equivalent figure in any Hurst work. As stated above, if the median and 95% CI are calculated for the 11 values of $\boldsymbol{{K}}$ in the in the [109, Table 1], there is an asymmetry between the medians and their 95% CI for the values of $K$, even though we do not place the values here because the samples become small, the asymmetry also exists for the sets of $\boldsymbol{{K}}$ separating the data of $\boldsymbol{{N=114}}$ and $N=25$ from [109, Table 1]. This probably ignorance of Hurst and the tendency until the first half of the twentieth century to consider all Gaussian data and call them “normal”, in the sense of being the “norm”.
@@
Hurst´s [109, Figure 2] shows that there is a linear relationship between $\boldsymbol{{log(\tfrac{R}{\sigma})}}$ and $\boldsymbol{{\log\left(N\right)}}$ for all sets in which there are enough groups. The equations of these lines are (for $\boldsymbol{{1.8\lessapprox\log(N)\lessapprox 3}}$ [109, in Figure 2])
$\log\left(\tfrac{R}{\sigma}\right)=K\log\left(\tfrac{N}{2}\right)=\log\left(R_{\sigma}\right).$ (37)
@@
$R_{\sigma}$ is called the rescaled range of fluctuations, in this case, of the Nile [160]. The value of $\boldsymbol{{K}}$ in Table 2 of Hurst [109] is positive and (calculated by me) 0.70(0.69 – $\boldsymbol{{0.70\;(0.69\text{ -- }0.72)}}$ (median and 95% confidence interval, calculated according to Hodges and Lehmann [100]), a small dispersion, perhaps somewhat skewed upwards. It is clear from the figures that for each set of phenomena a straight line fits the data well, and the remarkable fact that $\boldsymbol{{K}}$, the slope of the line, varies little from one set to another. A summary of the mean of the $K$ values obtained for the 690 separate values that Hurst computed, which is equivalent to $\boldsymbol{{\approx\tfrac{1,\bar{3}}{month}}}$ of the values that Hurst measured during years in which he devoted himself to that. The fraction $\boldsymbol{{\approx\tfrac{1,\bar{3}}{month}}}$ varies little, but probably reflects what would be expected for a large river: whose changes are slow. The number of data used $\boldsymbol{{\approx\tfrac{1}{month}}}$, is curious since some contemporary information suggests that the level of the Nile is measured daily [3].
@@
So far we have followed as closely as we could the data from Hurst [109]. Here we must point out, however, that Eq. (37) can be rewritten in general form as
@@
which suggests that straight lines are obtained by logarithms of any base. From here on we will join most authors calling $\boldsymbol{{K_{b}}}$ the Hurst’s coefficient and denote it as $\boldsymbol{{H}}$. An extremely important particular case is
$\begin{split}\ln\left(\dfrac{R}{\sigma}\right)&=\ln\left(R_{\sigma}\right)=H\ln\left(\dfrac{N}{2}\right)\\ R&=\left(\dfrac{N\sigma}{2}\right)\text{{{e}}}^{H}\implies R_{\sigma}=\left(\dfrac{N}{2}\right)\text{{{e}}}^{H}\\ \boldsymbol{{\text{\wasytherefore{\quad}}H}}&\boldsymbol{{=\ln{\left(\frac{2R}{N\sigma}\right)}\qquad\qquad\qquad\qquad\qquad\boldsymbol{\square}}}\end{split}$ (39)
If $R$ is Gaussian withh mean $\dot{\mu}$ and variance $\dot{\sigma}^{2}$ (that is $\boldsymbol{N}\left\{{\dot{\mu},\dot{\sigma}^{2}}\right\}$, as assumed by Hurst), it may also be written
$H=\ln\left(\frac{2}{\sigma N}\right)+\ln\left(\boldsymbol{N}\left\{{\dot{\mu},\dot{\sigma}^{2}}\right\}\right),$ (40)
so, even when $R$ is Gaussian, $H$ wont be.
Hurst [109, p. 19] considers the case where the annual water demand is less than the average river flow, but not less than the minimum annual flow recorded. In this case you have an approximate, $\boldsymbol{{B}}$, equal to what is required and less than the average, $\boldsymbol{{M}}$, and we need to know the largest accumulated deficit, $\boldsymbol{{S}}$, which must be covered with stored water. This can be determined for any particular record of the cumulative deviation curve.
For Figure 11, $\boldsymbol{{0\rightarrow X}}$ is the abscissa, which represents the value of the mean, with respect to which the other values of the figure are graphed. If we now produce an estimate $\boldsymbol{{B}}$ less than the average storage (axis $\boldsymbol{{0\rightarrow X}}$) it will be increased each year in the magnitude $\boldsymbol{{M-B}}$ above what was the estimated $\boldsymbol{{M}}$. This can be determined from the original curve of cumulative deviations by drawing an axis $\boldsymbol{{0\rightarrow X}}$ passing through the point $\boldsymbol{{F}}$ whose ordinate is $\boldsymbol{{-N(M-B)}}$. The curve ordinates referred to the $\boldsymbol{{0axis\rightarrow X_{1}}}$ show storage changes with water withdrawals $\boldsymbol{{B}}$. Storage drops from $\boldsymbol{{P}}$ to $\boldsymbol{{Q}}$, $\boldsymbol{{P\rightarrow c+d\rightarrow Q}}$. This is the amount of storage $S$ that would have been necessary to meet the demand $\boldsymbol{{B}}$.
$\boldsymbol{{S}}$ was determined for various water demands for each of $38$ types of phenomena such as the kinds of river discharges, rainfall, evaporation, use, and temperature that determines evaporation. The results for each class of events were grouped according to extraction, and the means of these groups were plotted. Evaporation was an important factor determining the salinization of Lake Nasser [6]. Two relationships fit to the results really well. The adjustment is (this strange equation is Hursy´s verbatim [109, Eq. (3)]), although the interrogation sign is mine:
@@
From there Hurst [109, Eq. (4)]) somehow gets
@@
The average value of $\boldsymbol{{N}}$ from which these results are calculated is $96$. It will be seen that in the adjustment, on the range of observations, there is no significant difference between one type of relationship and the other. If the figures are examined you will notice values of these kinds of phenomena are indistinguishable.
Hurst data on thr Nile river is indeed priceless, but as seen in Eqs. (41) and (42), his algebra is sometimes strange. Another source of uncertainty in Hurst´s data analysis is his unproven data Gaussianity, which is important to relate his exponent (now called H) with D, the fractal dimension of Nile´s water level fluctuations [143, 70, 71, Sutcliffe2016]. For me Hurst variable denomination is also strange and not easy to follow. All these peculiarities probably reflects Hurst’s isolation, and also perhaps his decision of starting to write on his work late in his life.
5.3 Validity of the relationship between the Hurst´s coefficient and its proposed relation with the fractal dimension.
The origin of the association between the two parameters was initially proposed by Mandelbrot and Wallis [146], with Mandelbrot’s success in finding fractal systems in the most diverse places, seemed to be evidence for a relationship between the Hurst exponent and the fractal dimension, $\Phi$, of two-dimensional systems was a simple relationship such as
@@
[139, 94, Sevcik1998a, Sevcik2010]. Thanks to Mandelbrot’s success, fractal analysis has been applied to time series, profiles, and natural surfaces in almost every scientific discipline.
More recent analyses [143, 70, 71, Sutcliffe2016] indicate that the relationship between fractal dimension and Hurst’s coefficient is actually more complex than initially thought, and is not described by Eq. (43).
The trouble with Hurst’s analysis is that he measured Nile River “levels” respect to some river bottom oe “zero” which is bot precised. By definition the Nile River level is never negative respect to this “zero”. Because of this, the measurements in Figure 11 are done to are some mean level which is never zero, but is taken as “zero”. And as indicated by my Eq. (39) (taken from Hurst work) $H$
Several authors have developed independent and (more?) exact methods to evaluate the Hurst coefficient [71, 68, Sanchez2015] which do not presuppose the Gaussianityd of the data that Hurst made and even consider the effect of the data ‘pathological’ distributed with Cauchy PDF [49, 169, Walck1996, Wolfram2003]. The discussion of “pathological distributions” like Cauchy’s is neither trivial nor irrelevant. These distributions are not Gaussian, and have neither mean nor variance. The data that Hurst discusses are deviations from the level of the Nile from its mean, estimated $\boldsymbol{{\bar{x}}}$, this is a random variable, whose mean is
@@
${\text{var}\left[{\overline{\Delta x_{i}}}\right]}=\dfrac{n+1}{n-1}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}.$ (45)
If $\boldsymbol{{x_{i}}}$ and $R$ [see Eq. (35)] were Gaussian as assumed by Hurst, then $\overline{\Delta x_{i}}$ and $R$ would have Cauchy PDF and no operation with its apparent mean and/or apparent variance would make sense. The Eqs. (44) and (45) can be calculated but are meaningless for a Cauchy distribution without central moments [79, 81, 82, 83, Theiler1990, 80]. If $\boldsymbol{{x_{i}}}$ is not Gaussian the problem is similar, an unknown pathological random variable whose mean and variance estimated from a sample, are meaningless, because the pdf of count come has no mean or defined variance. All sampling theory makes sense if and only if, the parameters that are estimated from the sample are approximations of the parameters of the population we sample.
@@
The idea that certain systems, numerically apparently very simple, could not have predictable values except at short times, was born when Lorenz [133] tried to solve a model of climate. The Lorenz model focused on three simultaneous equations (with slightly modified notation) which ate a numerical approximation to Rquation System (1):